Can I please get help with:

(12)/(3x-2)+(5)/(x+4)=1

Thanks you.

Answer by edjones(8007)

(12)/(3x-2)+(5)/(x+4)=1
12(x+4)+5(3x-2)=(3x-2)(x+4) multiply each side by the LCM to eliminate fractions.
12x+48+15x-10=3x^2+10x-8
27x+38=3x^2+10x-8
3x^2-17x-46=0 Subtract the left side from the right.
3*-46=-138
factors of -138 that add up to -17: 6 & -23
3x^2+6x-23x-46=0
3x(x+2)-23(x+2)=0 Factoring by grouping. (easier)
(3x-23)(x+2)=0
x=-2, x=23/3
.
Ed